### v1

# Adaptive Testing for LLM Evaluation: A Psychometric Alternative to Static Benchmarks

Abstract
Large language model evaluation requires thousands of benchmark items, making evaluations expensive and slow. Existing methods compute average accuracy across fixed item sets, treating all items equally despite varying quality and informativeness. We present ATLAS  an adaptive testing framework using Item Response Theory (IRT) to estimate model ability through Fisher information-guided item selection. Our analysis of five major benchmarks reveals that 3-6% of items exhibit negative discrimination, indicating annotation errors that corrupt static evaluation. ATLAS achieves 90% item reduction while maintaining measurement precision: on HellaSwag (5,608 items), we match full-benchmark estimates using only 42 items with 0.154 MAE. Our framework maintains item exposure rates below 10% and test overlap at 16-27%, compared to static benchmarks where every model sees all items (100% exposure). Among 4,000+ tested models, IRT ranks differ from accuracy ranks: models with the same accuracy get different IRT scores, and 23-31% of all models shift by more than 10 rank positions. Code and calibrated item banks are available at https://github.com/Peiyu-Georgia-Li/ATLAS.git.
**footnotetext: These authors contributed equally.
1 Introduction
Large language model (LLM) evaluation relies on benchmarks with tens of thousands of test items. These benchmarks impose high computational costs and evaluation cycles spanning days or weeks. Despite benchmark growth—some proposals exceed 100,000 items—evaluation practice remains static: models receive average accuracy scores across fixed item sets. This approach ignores statistical information in large datasets and raises questions about efficiency and validity.
Current evaluation faces three fundamental limitations. First, average scores hide meaningful differences between models with distinct error patterns, particularly for lower-performing models where small ability differences are obscured by measurement noise. Second, static benchmarks create vulnerability to data contamination as items leak into pretraining corpora, enabling high scores through memorization rather than genuine capability. Third, evaluating complete benchmarks is inefficient, treating poorly discriminative items as equally informative as high-quality questions.
To address these limitations, we propose ATLAS (Adaptive Testing for LLM Ability Scoring), an adaptive evaluation framework based on computerized adaptive testing (CAT; (lord1980, ; wainer2000computerized, ; weiss1982improving, )). ATLAS first calibrates benchmark items using three-parameter logistic (3PL) IRT models to estimate item difficulty, discrimination, and guessing parameters (birnbaum1968some, ; hambleton1991fundamentals, ). Then, rather than administering fixed item sets, ATLAS dynamically selects items with maximum Fisher information for each model’s current estimated ability, terminating when precision thresholds are reached. This approach directly addresses all three limitations: Fisher information-guided selection provides precise ability estimates that distinguish models with identical accuracy, dynamic item selection reduces contamination through diverse test forms, and adaptive termination achieves reliable evaluation with significantly fewer items.
We validate our framework across five major benchmarks (WinoGrande, TruthfulQA, HellaSwag, GSM8K, ARC). ATLAS achieves competitive or superior performance to static baselines: on TruthfulQA, it attains the lowest MAE (0.067) using only 51 items; on HellaSwag, it matches MetaBench accuracy (0.154 MAE) while requiring 3× fewer items (39 vs 93). Across benchmarks, ATLAS requires 30-78 items on average compared to thousands in full evaluation, while IRT-based estimates reveal systematic rank reordering that exposes limitations of accuracy-based evaluation.
Our contributions are: (1) We demonstrate fundamental limitations of average-score evaluation and establish the benefits of psychometric ability estimates for LLM evaluation. (2) We develop ATLAS, the first large-scale adaptive testing framework for LLMs, achieving 90% item reduction while maintaining measurement precision and natural contamination resistance through $10\%$ item exposure rates. (3) We provide comprehensive psychometric analysis of five major benchmarks, revealing that IRT-based estimates cause systematic rank reordering ($23-31\%$ of models shift more than 10 positions) and identifying that $3-6\%$ of items exhibit negative discrimination due to annotation errors, with calibrated item banks released for reproducible evaluation.
2 Related Work
2.1 Background
Large-scale benchmarks such as WinoGrande, TruthfulQA, HellaSwag, GSM8K and ARC have become standard for LLM evaluation but impose significant computational costs. These static approaches suffer from benchmark saturation, data contamination, and item redundancy, raising concerns about their ability to reflect genuine model capabilities.
Item Response Theory (IRT) has recently been applied to LLM evaluation (lalor2024item, ; guinet2025evaluating, ). It provides item parameters such as difficulty, discrimination, and guessing, as well as latent ability estimates $\theta$ for models. However, existing IRT applications remain largely static: tinyBenchmarks (polo2024tinybenchmarks, ) uses clustering for item selection but doesn’t guarantee informativeness for $\theta$ estimation, while MetaBench (metabench, ) requires computationally expensive iterations to identify stable subsets. These approaches often lack proper psychometric validation and emphasize predictive accuracy over model fit diagnostics, which makes it difficult to ensure that ability estimates are valid, interpretable, and comparable across models.
Beyond these limitations of existing IRT applications, many evaluations continue to rely on average scores. Average scores tend to mask meaningful model differences and are often affected by form-dependence, nonlinear scaling, equal weighting of uninformative items, and contamination sensitivity (see Appendix A for detailed analysis). In contrast, IRT-based ability estimates ($\theta$) provide form-invariant, uncertainty-aware alternatives that adjust for item difficulty and discrimination.
2.2 Adaptive Testing
Computerized adaptive testing (CAT) is a psychometric framework for efficiently measuring ability by selecting items that maximize Fisher information. By adapting item selection in real time, CAT caters to different ability levels, administering easier items to weaker models and more challenging items to stronger ones. CAT can operate with a fixed test length or with a variable length, terminating once a predefined precision threshold is reached (weiss1982improving, ).
The challenge of evaluating LLMs closely parallels that of international large-scale assessments for humans. Our dataset includes more than 4,000 models, ranging from simple pattern-matching systems with very low ability estimates ($\theta<-2$) to advanced models such as GPT-4 with high ability estimates ($\theta>2$) . This wide heterogeneity makes static benchmarks inadequate: fixed test forms inevitably include items that are trivial for advanced models and impossible for weaker ones, which limits their diagnostic value. Unlike static benchmarks, CAT explicitly adapts to each model’s ability level, ensuring that item selection remains informative across the full spectrum of performance.
CAT allows us to reduce test length while maintaining measurement precision. Modern extensions such as Multi-Stage Testing and the incorporation of process data (zenisky2009multistage, ; zheng2015fly, ; tang2024utilizing, ) provide further gains in efficiency and validity. Prior attempts to explore adaptive evaluation for LLMs have been either limited in scope (zhuang2023efficiently, ) or primarily conceptual (zhuangposition, ). Beyond efficiency, CAT also strengthens defenses against contamination. Because models are evaluated on small, adaptively chosen subsets selected in real time, the overlap between training and evaluation data is substantially reduced. Furthermore, CAT protocols can be refreshed with new items while preserving comparability of ability estimates $\theta$. A detailed comparison of IRT-based approaches is provided in Appendix B.
3 Methodology
We introduce a novel adaptive testing framework that transforms LLM evaluation from static benchmarking to dynamic ability estimation. Our approach addresses three critical limitations of current evaluation practice: (1) it distinguishes models with identical average scores but different capability patterns, (2) it reduces computational cost by requiring 90% fewer items while maintaining accuracy, and (3) it naturally mitigates contamination risks through diverse item sampling.
This section presents our framework in four stages: problem formulation (Section 3.1), data construction with psychometric filtering (Section 3.2), item bank calibration using IRT models (Section 3.3), and adaptive testing with randomesque selection (Section 3.4).
3.1 Problem Formulation and Setup
We formulate LLM evaluation as a psychometric measurement problem. Let $\mathcal{I}$ denote the set of benchmark items and $\mathcal{L}$ the set of language models. For each model $\ell\in\mathcal{L}$ and item $i\in\mathcal{I}$, we observe a binary response $Y_{i,\ell}\in\{0,1\}$, where 1 indicates correct and 0 incorrect. These responses form the item-response matrix $\{Y_{i,\ell}\}_{i\in\mathcal{I},\ell\in\mathcal{L}}$.
Unlike traditional approaches that rely solely on accuracy scores, our objective is to estimate the latent ability $\theta_{\ell}$ of each model based on its response pattern $\{Y_{i,\ell}\}_{i\in\mathcal{I}}$, while simultaneously calibrating item-level parameters: discrimination $a_{i}$, difficulty $b_{i}$, and guessing $c_{i}$. This approach enables fine-grained model comparison even when models achieve identical accuracy, as $\theta_{\ell}$ accounts for the varying informativeness of different items.
3.2 Discriminative Item Filtering
We construct the item-response matrix using data from the HuggingFace Open LLM Leaderboard. The item pool $\mathcal{I}$ spans five benchmarks: ARC, GSM8K, HellaSwag, TruthfulQA, and WinoGrande. To ensure data quality for IRT calibration, we apply two levels of filtering: removing unsuitable models and eliminating non-informative items.
Model Selection. We retain only models $\mathcal{L}$ with complete responses across all items to ensure unbiased ability estimates. Models with extreme scores (below 0.1st percentile) are excluded to prevent parameter estimation instability, as IRT’s sigmoidal functions become under-constrained at the boundaries.
Item Filtering. We apply two complementary filters to retain only discriminative items:
  • Low-variance removal: Items with response standard deviation $1\%$ or mean accuracy $95\%$ are discarded, as they provide litter information for differentiating between models.
  • Discrimination filtering: We compute the point-biserial correlation $r_{pb}(i)$ between each item’s response vector $\{Y_{i,\ell}\}_{\ell\in\mathcal{L}}$ and the models’ total scores $T_{\ell}=\sum_{j\in\mathcal{I}}Y_{j,\ell}$ (see Appendix E for details). Items with $r_{pb}(i)<0.1$ are removed as non-diagnostic.
This filtering process yields a refined response matrix that supports stable and reliable IRT calibration (see Table 5 in Appendix E for detailed results).
3.3 Scalable IRT Calibration
The calibration stage estimates item parameters $(a_{i},b_{i},c_{i})$ and computes reference ability estimates $\hat{\theta}_{\ell}^{\mathrm{whole}}$ for each LLM $\ell$ for validation. To model the probability of a correct response, we adopt the three-parameter logistic (3PL) IRT model (birnbaum1968some, ; lord1980, ):
$p_{i}(\theta_{\ell})=c_{i}+\frac{1-c_{i}}{1+\exp(-a_{i}(\theta_{\ell}-b_{i}))}.$ (1)
Here, $a_{i}$ is the discrimination parameter, which determines how sharply item $i$ differentiates between stronger and weaker models. $b_{i}$ is the difficulty parameter, specifying the ability level at which a model has a 50% chance (beyond guessing) of answering item $i$ correctly. $c_{i}$ is the guessing parameter, setting the lower bound on the probability of success due to random guessing. These parameters enable Fisher information-based prioritization of items in our adaptive framework, distinguishing high-quality items from those with low discriminative power.
Common-Person Calibration at Scale. We develop a partition-based calibration strategy that takes advantage of the unique characteristics of LLM benchmarking. Unlike traditional testing where only subsets of items are administered to each examinee, all LLM models respond to all items, enabling a novel application of common-person linking. We partition items into $K$ non-overlapping subsets $\mathcal{I}_{k}$ (each with $|\mathcal{I}_{k}|\geq 100$ items), calibrate each subset independently, then apply mean-sigma transformations (kolen2014, ) using the model population $\mathcal{L}$ as linking anchors. This approach reduces computational complexity from $O(|\mathcal{I}|^{3})$ to $O(K\cdot\max_{k}|\mathcal{I}_{k}|^{3})$ while maintaining calibration accuracy through the redundancy of having all models as linkers (chalmers2012mirt, ).
Heterogeneity-Aware Ability Estimation. The extreme heterogeneity of LLM populations, spanning from random baselines ($\theta\approx-3$) to near-perfect models ($\theta\approx 3$), breaks standard estimation methods. Standard IRT practice relies on Expected A Posteriori (EAP) estimation (bock1982adaptive, ) for ability estimation. However, EAP produces undefined values for all-correct or all-incorrect response patterns, affecting 12% of our model population spanning from random baselines to near-perfect systems. To address this LLM-specific challenge, we employ Weighted Likelihood Estimation (WLE) (warm1989weighted, ) with bias correction term $\frac{J(\theta)}{2I(\theta)}$ where $J(\theta)=\sum_{i}\frac{\partial I_{i}(\theta)}{\partial\theta}$ (warm1989weighted, ). This choice enables finite estimates across the full ability spectrum while maintaining consistency properties (baker2004item, ), crucial for establishing reliable validation baselines in the extreme heterogeneity of LLM evaluation.
Multi-Subset Model Fit Validation. Unlike prior IRT applications to LLM evaluation (polo2024tinybenchmarks, ; metabench, ) that omit model validation, we conduct rigorous psychometric diagnostics to ensure calibration quality. We report the $M_{2}$ statistic (maydeu2015m2, ) with RMSEA indices, establishing that our 3PL models achieve good fit (RMSEA <0.05<0.05) across all benchmarks, as summarized in Appendix D. Our partition-based calibration strategy, which divides the full item bank into non-overlapping subsets (each containing $\geq 100$ items for statistical stability), enables both computational feasibility and robust validation. Since the same set of models $\mathcal{L}$ acts as common persons across all partitions, diagnostic statistics reflect global calibration quality rather than partition-specific artifacts. This multi-subset linking design ensures that model fit metrics capture systematic patterns across the entire item bank, not just localized subsets. This validation is crucial for reliable adaptive testing, as misfitting items would compromise Fisher information calculations and degrade selection accuracy.
3.4 Adaptive Testing with Information Selection
Our proposed ATLAS dynamically selects the most informative items for each model, dramatically reducing the number of items needed while maintaining accuracy. Algorithm 1 presents the complete procedure. The algorithm includes several key design choices tailored to LLM evaluation:
Algorithm 1 Adaptive Testing for Model $\ell$
1:Initialize: $\hat{\theta}_{0}\leftarrow 0$, test record $R_{\ell}\leftarrow\emptyset$, $t\leftarrow 0$
2:while $t<\mathrm{max\_items}$ and not converged do
3:  $t\leftarrow t+1$
4:  if $t=1$ then
5:   Select item $i_{t}$ with $|b_{i_{t}}-\hat{\theta}_{0}|$ minimized
6:  else
7:   Compute Fisher information $I_{i}(\hat{\theta}_{t-1})$ for all unadministered items
8:   Select $i_{t}$ randomly from top-5 most informative items
9:  end if
10:  Administer item $i_{t}$ to model $\ell$, observe response $Y_{i_{t},\ell}$
11:  Update record: $R_{\ell}\leftarrow R_{\ell}\cup\{(i_{t},Y_{i_{t},\ell})\}$
12:  Update ability: $\hat{\theta}_{t}\leftarrow\mathrm{EAP}(R_{\ell})$
13:  Compute standard error: $\mathrm{SE}(\hat{\theta}_{t})\leftarrow 1/\sqrt{\sum_{j\in R_{\ell}}I_{j}(\hat{\theta}_{t})}$
14:  if $t\geq\mathrm{min\_items}$ and $\mathrm{SE}(\hat{\theta}_{t})\leq\tau$ then
15:   break $\triangleright$ Convergence achieved
16:  end if
17:end while
[…]
Initialization and Bounds. We initialize the ability estimate at $\hat{\theta}_{0}=0$. In IRT, abilities are typically standardized to mean zero, so this choice provides a neutral starting point that avoids bias toward high- or low-performing models and stabilizes early item selection when little information is available. We enforce minimum (30) and maximum (500) item limits. The minimum ensures stable estimation for models at performance extremes, while the maximum constrains computational cost and yields approximately 90% reduction in test length relative to full benchmarks.
Randomesque Item Selection. Rather than deterministically selecting the single most informative item, we randomly sample from the top-5 candidates ranked by Fisher information:
$I_{i}(\theta)=a_{i}^{2}\cdot p_{i}(\theta)\cdot[1-p_{i}(\theta)].$ (2)
This randomesque strategy (kingsbury1989procedures, ) prevents over-reliance on specific item types while still keeping high information, which is important for models with specialized capabilities.
Sequential Ability Updates. After each item administration, we update the ability estimate using Expected A Posteriori (EAP) estimation (bock1982adaptive, ):
$\hat{\theta}_{t}=\mathbb{E}[\theta|R_{\ell}]=\int\theta\cdot p(\theta|R_{\ell})\,d\theta.$ (3)
EAP provides numerically stable updates with sparse early responses and incorporates prior knowledge about ability distributions. In contrast, WLE tends to become unstable when response patterns are extreme, a situation common in the early stages of adaptive testing.
[…]
Output and Validation. For each model $\ell$, the algorithm produces: (1) the administered item sequence and responses $R_{\ell}$, (2) the ability estimate trajectory $\{\hat{\theta}_{t}\}$ with associated standard errors, and (3) the final estimate $\hat{\theta}_{\ell}$. We validate these adaptive estimates against whole-bank references $\hat{\theta}_{\ell}^{\mathrm{whole}}$ to confirm that our dramatic reduction in items does not compromise measurement accuracy.
4 Experiments
We evaluate the proposed ATLAS framework across five benchmarks, comparing its efficiency and accuracy against static baselines while analyzing exposure rates, overlap patterns, and the advantages of IRT-based ability estimation over raw accuracy scores.
4.1 Experimental Setup and Metrics
We evaluate ATLAS across five diverse benchmarks covering different cognitive domains: WinoGrande (commonsense reasoning), TruthfulQA (factual consistency), HellaSwag (procedural inference), GSM8K (mathematical reasoning), and ARC (scientific question answering). All experiments use calibrated item banks from Section 3.3.
We compare against four static baselines that do not adapt to individual models: (1) Random sampling of 100 items uniformly from the full bank, (2) TinyBenchmarks (polo2024tinybenchmarks, ) using predetermined subsets selected via clustering without Fisher information optimization, (3) MetaBench-Primary and (4) MetaBench-Secondary (metabench, ) using curated splits that require computationally expensive iterations to identify stable subsets. Unlike these static approaches, ATLAS uses three precision thresholds ($\mathrm{SE}(\hat{\theta})\leq 0.1,0.2,0.3$) with bounds of 30–500 items, terminating when sufficient precision is achieved.
We evaluate using five metrics: (1) Mean Absolute Error (MAE) between ATLAS estimates $\hat{\theta}_{\ell}$ and full-bank references $\hat{\theta}_{\ell}^{\text{whole}}$ to assess accuracy, (2) Average Number of Items Administered to measure efficiency, (3) Item Exposure Rate quantifying how frequently each item appears across evaluations, (4) Test Overlap Rate $\bar{Q}$ measuring the average proportion of shared items between any two model evaluations, and (5) Runtime per session to assess computational efficiency. The exposure rate is calculated as the percentage of models that encounter each specific item, while the overlap rate represents the average Jaccard similarity between item sets administered to different models. Additional experimental configurations and psychometric interpretations are provided in Appendix C.
Table 1: Baseline and adaptive evaluation results (MAE and Avg. Item). Lower is better $\downarrow$.
Method WinoGrande TruthfulQA HellaSwag GSM8K ARC
MAE Items MAE Items MAE Items MAE Items MAE Items
Random 0.176 100 0.105 100 0.231 100 0.164 100 0.147 100
TinyBenchmarks 0.247 100 0.161 100 0.299 100 0.173 100 0.165 100
MetaBench-P 0.148 133 0.081 154 0.164 93 0.154 100 0.142 145
MetaBench-S 0.210 106 0.074 136 0.238 58 0.162 100 0.133 100
ATLAS0.1 0.149 78 0.067 51 0.154 39 0.159 73 0.099 77
ATLAS0.2 0.181 39 0.074 30 0.159 30 0.181 39 0.126 36
ATLAS0.3 0.183 32 0.072 30 0.161 30 0.186 32 0.128 30
Table 2: Adaptive evaluation efficiency and diversity. ATLAS maintains low item exposure rates ($10\%$) and moderate test overlap ($16-27\%$), providing natural contamination resistance while achieving fast selection times (<1<1 minute per model). Lower values indicate better performance for all metrics.
Benchmark Method Test Overlap $\downarrow$ Avg. Item $\downarrow$ Avg. Selection $\downarrow$
(Item #) Rate (%) Exposure Rate (%) Time (s)
WinoGrande
ATLASSE≤0.1 18.85 9.28 40.99
ATLASSE≤0.2 17.32 5.07 19.92
ATLASSE≤0.3 20.28 4.24 16.74
TruthfulQA
ATLASSE≤0.1 18.93 8.17 15.97
ATLASSE≤0.2 21.69 10.56 9.37
ATLASSE≤0.3 21.67 10.68 9.72
HellaSwag
ATLASSE≤0.1 15.76 3.90 75.52
ATLASSE≤0.2 19.28 6.23 56.93
ATLASSE≤0.3 19.25 6.85 57.06
GSM8K
ATLASSE≤0.1 22.43 8.32 45.69
ATLASSE≤0.2 22.16 5.63 24.08
ATLASSE≤0.3 26.89 8.24 19.06
ARC
ATLASSE≤0.1 19.50 10.08 30.99
ATLASSE≤0.2 19.25 5.42 13.98
ATLASSE≤0.3 21.91 9.77 11.82
4.2 Performance and Reliability Analysis
Figure 1: Comparison of subset (predicted) ability estimates against whole-bank (reference) abilities across five benchmarks. Points along the diagonal indicate perfect agreement. ATLAS maintains the closest alignment overall, particularly on TruthfulQA, ARC, HellaSwag and in the low-ability regime of GSM8K, while static baselines such as TinyBenchmarks and random show greater variance and systematic deviation.
Table 1 reveals ATLAS’s consistent efficiency advantages across all benchmarks. On HellaSwag, ATLAS achieves the best accuracy (0.154 MAE) using only 39 items - just 42% of MetaBench-Primary’s 93 items while improving accuracy by 6%. The efficiency gains are most dramatic on TruthfulQA, where ATLAS achieves the lowest MAE of 0.067 using only 33% of MetaBench-Primary’s items (51 vs 154) and outperforms both MetaBench variants. On ARC, ATLAS demonstrates substantial improvements in both dimensions: 30% better accuracy (0.099 vs 0.142 MAE) using just 53% of the items (77 vs 145). Even on GSM8K where MetaBench-Primary achieves marginally lower error, ATLAS maintains competitive performance (0.159 vs 0.154 MAE) while requiring 27% fewer items. WinoGrande shows similar patterns, with ATLAS matching MetaBench-Primary’s accuracy (0.149 vs 0.148 MAE) using only 59% of the items.
Figure 1 compares ability estimates from subsets against whole-bank references. ATLAS’s tight alignment to the diagonal demonstrates reliable ability estimation across benchmarks. In contrast, static baselines show systematic deviations: TinyBenchmarks exhibits bias at high-ability levels, while random sampling produces broad scatter, confirming the importance of principled item selection over arbitrary subsampling.
4.3 Contamination Resistance and Efficiency Benefits
Table 2 demonstrates ATLAS’s contamination resistance and efficiency advantages across multiple dimensions. The average item exposure rates remain below 10% across all benchmarks and precision levels, with the lowest rates on large banks: HellaSwag achieves just 3.9% exposure with SE $\leq 0.1$, meaning each item appears in fewer than 4% of model evaluations. This contrasts sharply with static baselines where every item is exposed in 100% of evaluations. Even smaller banks maintain low exposure rates, with ARC and TruthfulQA showing 5.4-10.7% exposure rates, making systematic memorization during pretraining practically impossible.
The test overlap rates of 16-27% provide natural contamination resistance while maintaining evaluation consistency. On HellaSwag, the 15.8% overlap rate means two randomly selected models share fewer than 16% of items on average, compared to 100% overlap in static evaluation. This diversity scales with bank size: larger banks like HellaSwag (5,608 items) achieve the lowest overlap rates (15.8-19.3%), while smaller banks like GSM8K show higher but still protective overlap rates (22.2-26.9%). The precision threshold affects this trade-off: stricter thresholds (SE $\leq 0.1$) generally produce lower overlap rates but higher exposure rates due to longer tests.
Runtime efficiency demonstrates practical scalability, with selection times averaging 9.4-75.5 seconds per model across benchmarks. The timing scales roughly with bank size: HellaSwag requires the longest selection time (57-76 seconds) due to its large item pool, while TruthfulQA achieves the fastest selection (9.4-16.0 seconds). Importantly, these times include the full adaptive selection process and remain well under one minute for most configurations, making ATLAS practical for both interactive evaluation and large-scale benchmarking scenarios.
(a) GSM8K
(b) HellaSwag
Figure 2: Comparison of IRT ability estimates $\hat{\theta}_{\ell}^{\text{whole}}$ with raw accuracy. Left: Ability vs accuracy reveals strong correlation but critical differences at performance extremes where accuracy collapses. Right: Rank comparison shows systematic reordering, with 23% (GSM8K) and 31% (HellaSwag) of models shifting >> 10 positions. IRT separates models with identical accuracies by accounting for which items they solve correctly.
4.4 Distinguishing Low- and High-Performing Models
Figure 2 demonstrates IRT’s key advantage: separating models with similar accuracy scores through ability estimates. Despite strong correlations (0.99 for GSM8K, 0.96 for HellaSwag), systematic differences emerge at performance extremes. In low-performing regimes, accuracy collapses into narrow bands (0.10-0.15) while IRT spans $\theta\approx-3$ to 1-1, distinguishing models that solve easy versus challenging items. In high-performing regimes, ceiling effects compress accuracy differences, but IRT maintains discrimination across $\theta\approx 1.5$ to $2.5$. Similar patterns are observed across TruthfulQA, WinoGrande, and ARC benchmarks, with additional score-versus-theta comparisons provided in Appendix G.
The right panels show systematic rank reordering: 23-31% of models shift >> 10 positions when ranked by IRT versus accuracy. Models performing well on hard items receive higher IRT ranks despite moderate accuracy, while those succeeding on easy items are appropriately downweighted. This enhanced discrimination provides more reliable model comparisons, especially critical in saturated performance regions where accuracy-based evaluation fails. Similar patterns of IRT superiority in distinguishing models are observed across TruthfulQA, WinoGrande, and ARC benchmarks, with additional analysis provided in Appendix G.
4.5 Items Are Not Equally Informative
Figure 3: Two models with identical accuracy (0.833) on WinoGrande receive different ability estimates ($\hat{\theta}_{A}=1.2$ vs $\hat{\theta}_{B}=0.6$). Model A succeeds on harder items (darker cells on right), while Model B answers easier items (darker cells on left). IRT captures these item difficulty patterns that raw accuracy cannot.
Figure 3 demonstrates IRT’s key advantage: models with identical accuracy (0.833) receive different ability estimates ($\hat{\theta}_{A}=1.2$ vs $\hat{\theta}_{B}=0.6$) based on which items they solve correctly. Under the 3PL IRT model, items differ in discrimination $a_{i}$, difficulty $b_{i}$, and guessing $c_{i}$ parameters. Model A succeeds on high-difficulty items ($b_{i}>0.5$) with strong discrimination ($a_{i}>1.5$), yielding 2.3× more Fisher information than Model B, which mainly answers easier, less discriminative items ($b_{i}<-0.5$, $a_{i}<0.8$).
IRT provides automatic quality control by weighting items according to their empirical contribution to distinguishing model abilities. In our calibrated banks, 3.2-5.7% of items exhibit negative discrimination ($a_{i}<0$), indicating systematic flaws where stronger models perform worse. For example, WinoGrande item #247 achieves 0.89 accuracy but $a_{i}=-0.43$ due to exploitable linguistic artifacts. Under raw accuracy, this flawed item contributes equally (weight = 1/N) to all scores, potentially inflating weak models. Under IRT, negative discrimination automatically down-weights its contribution by 82% (effective weight $\propto a_{i}^{2}\approx 0.18$), reducing measurement contamination and providing more reliable ability estimates than accuracy alone. Similar patterns of identical accuracy leading to different ability estimates are observed across ARC, HellaSwag, and other benchmarks, with additional heatmap visualizations provided in Appendix G.
5 Discussion
5.1 Implications for Benchmark Design
Our findings underscore several critical considerations for the design of future large language model (LLM) benchmarks. First, reduced item sets necessitate content balancing to preserve assessment validity. Content balancing entails maintaining proportional representation across distinct skill domains or cognitive competencies to ensure that the benchmark reflects the intended construct rather than overemphasizing specific subskills. Without such balance, benchmarks risk producing biased or unrepresentative evaluations. The proposed adaptive framework can be readily extended to optimize both domain coverage and measurement precision simultaneously (cheng2009maximum, ). Achieving comparable balance using static, reduced item sets would be substantially more difficult and would require extensive manual tuning or domain expertise.
Second, item quality exerts a direct influence on evaluation reliability. Our analysis reveals that 3.2–5.7% of items display negative discrimination parameters, indicating systematic flaws that undermine accuracy-based performance estimates. Although IRT automatically attenuates the influence of such items through probabilistic weighting, many existing benchmarks lack analogous quality control mechanisms. Consequently, poorly constructed or ambiguous items can distort aggregate scores and obscure genuine model differences. Future benchmark design should therefore adopt psychometric validation as standard practice, incorporating procedures such as discrimination screening and content alignment checks to enhance the interpretability and fairness of evaluation outcomes.
Third, item exposure rate and test overlap rate are essential metrics for benchmark robustness and reusability. A low exposure rate ensures that individual items are rarely reused across assessments, reducing the risk of model memorization or contamination of test content. Similarly, a low test overlap rate—the proportion of shared items between distinct benchmark forms—preserves evaluation independence and minimizes redundancy. Adaptive benchmarking intrinsically achieves lower exposure and overlap rates by tailoring item selection to each model’s estimated proficiency, generating diverse yet psychometrically comparable test forms. This adaptivity confers a distinct advantage over static benchmarks, enabling more secure, scalable, and sustainable evaluation practices in iterative model development contexts.
5.2 Limitations and Future Work
Despite notable efficiency gains, our framework has several limitations. Initial calibration relies on a representative model population, which may become outdated as architectures evolve. Nonetheless, adaptive testing supports incremental updates: new model responses can be incorporated to refine item parameters and maintain calibration accuracy over time. The current implementation remains limited to multiple-choice formats, while generative or open-ended tasks require alternative scoring and modeling strategies. Moreover, our framework is unidimensional, assuming a single latent proficiency dimension across all items, whereas LLM capabilities are inherently multidimensional. Future work should extend the framework to multidimensional IRT formulations that jointly model reasoning, factuality, and linguistic ability.
To facilitate broader adoption, we have implemented a modular scoring interface that allows users to define custom evaluation functions (e.g., determining whether a model answers a selected item correctly). Future work will explore online or Bayesian calibration techniques for continuous item updating, multidimensional modeling to capture diverse LLM capabilities, and hybrid adaptive–generative evaluation for open-ended tasks. We also plan to extend the system with interfaces for cross-model benchmarking, enhanced item exposure control, and visualization tools that improve interpretability and diagnostic insight.
6 Conclusion
Adaptive testing replaces exhaustive benchmarking with efficient, contamination-resistant measurement. Our ATLAS framework reduces items by 90% while maintaining precision, separates models with identical accuracies using IRT-based ability estimates, and limits contamination through << 10% item exposure rates. With benchmark saturation and contamination risks increasing, adaptive testing provides a rigorous alternative, applying measurement science principles instead of static averaging.
Appendix A Detailed Analysis of Average Score Limitations
Average score (percent correct) remains the most widely reported metric for evaluating LLMs. While it provides a convenient ordinal indicator for fixed forms, it is a shaky measure of underlying ability.
First, average scores are form-dependent: changing the mix or difficulty of items alters percent correct, even if the model’s true ability is unchanged. Second, the metric has a nonlinear scale: improvements at the extremes (e.g., $98\%\!\to 100\%$) do not reflect the same underlying gain as improvements in the middle (e.g., $50\%\!\to 52\%$). Third, it assumes equal informativeness across items, allowing easy or guessable items to influence the mean as much as highly discriminative ones. Fourth, it is subject to coverage bias: the observed score reflects the content blueprint of the test rather than ability across domains. Fifth, average scores offer no measure of uncertainty, making it unclear whether differences are statistically meaningful. Finally, they are highly sensitive to contamination: memorized items from pretraining can artificially inflate percent correct without reflecting genuine reasoning or generalization.
In contrast, IRT-based ability estimates ($\theta$) provide form-invariant, uncertainty-aware measures that adjust for item difficulty and discrimination. Reporting $\theta\pm\text{SE}(\theta)$ offers a psychometrically principled alternative. For communication purposes, reconstructed percent scores may be shown alongside, but $\theta$ should serve as the primary indicator of model capability.
Appendix B Comparison of IRT-Based Benchmark Methods
Table 3: Comparison of IRT-Based Benchmark Methods for LLMs
Factor TinyBenchmarks (Static) MetaBench (Static) ATLAS
IRT Calibration Required Required Required
Adaptivity None (same items) None (same items) High (items vary by ability)
Test Length Fixed Fixed Variable, stopping rules
Exposure control High (same items reused) High (same items reused) Low (rotating pool)
Pool Sensitivity Subset dependent Subset dependent Robust to large pools
Fairness Biased if mis-targeted Biased if mis-targeted Balanced across abilities
Score Precision Low at extremes Low at extremes High, SEs available
Model Fit Rarely checked Rarely checked Possible fit checks
Efficiency Fixed (100 items) Fixed (100 items) Fewer items, comparable accuracy
Saturation Risk High High Low
Appendix C Detailed Experimental Setup and Metrics
C.1 Benchmarks and Datasets
We conduct experiments on five diverse benchmarks covering different cognitive domains:
  • WinoGrande: Commonsense reasoning with pronoun resolution
  • TruthfulQA: Factual consistency and truthfulness evaluation
  • HellaSwag: Procedural inference and common sense completion
  • GSM8K: Mathematical word problems requiring multi-step reasoning
  • ARC: Scientific question answering across multiple domains
All experiments use the calibrated item banks from Section 3.3, ensuring consistent filtering and parameter quality across datasets. Items with poor psychometric properties (negative discrimination $a_{i}<0$, extreme difficulty $|b_{i}|>4$, or unrealistic guessing $c_{i}>0.5$) were excluded during calibration.
C.2 Baseline Configurations
We compare ATLAS against four fixed, non-adaptive strategies:
Random Baseline: Samples 100 items uniformly from the full bank without consideration of item parameters or model ability.
TinyBenchmarks: Uses the predetermined subset from [8], selected via clustering methods but without explicit Fisher information optimization for ability estimation.
MetaBench-Primary and MetaBench-Secondary: Curated splits from [9] that require computationally expensive iterations to identify stable subsets. These splits emphasize predictive accuracy over psychometric validity.
All baseline data is available on Hugging Face: tinyBenchmarks, HCAI/metabench.
Unlike ATLAS, these approaches do not adapt to individual test-takers and serve only as static reference points for accuracy–efficiency tradeoffs.
C.3 ATLAS Configuration Details
For each model $\ell$, we run ATLAS under three precision-based stopping thresholds:
  • $\mathrm{SE}(\hat{\theta})\leq 0.1$: High precision, suitable for fine-grained model comparison
  • $\mathrm{SE}(\hat{\theta})\leq 0.2$: Moderate precision, balancing accuracy and efficiency
  • $\mathrm{SE}(\hat{\theta})\leq 0.3$: Lower precision, maximizing efficiency for rapid screening
A minimum of 30 items is enforced to prevent premature termination due to lucky guesses or initial high-information items, while the maximum is capped at 500 items to ensure computational feasibility. This setup balances precision and budget constraints, simulating realistic conditions for adaptive evaluation in production environments.
C.4 Detailed Metric Definitions
Average Mean Absolute Error (MAE):
$\text{MAE}=\frac{1}{|\mathcal{M}|}\sum_{\ell\in\mathcal{M}}|\hat{\theta}_{\ell}-\hat{\theta}_{\ell}^{\text{whole}}|$
where $\mathcal{M}$ is the set of evaluated models, $\hat{\theta}_{\ell}$ is the ATLAS-derived ability estimate, and $\hat{\theta}_{\ell}^{\text{whole}}$ is the reference estimate from the complete item bank.
Average Item Exposure Rate:
$\text{Exposure}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\frac{\text{count}(i)}{|\mathcal{M}|}$
where $\text{count}(i)$ is the number of times item $i$ was administered across all models $\mathcal{M}$, and $|\mathcal{I}|$ is the total number of items in the bank.
Test Overlap Rate:
$\bar{Q}=\frac{1}{\binom{|\mathcal{M}|}{2}}\sum_{\ell_{1}<\ell_{2}}\frac{|\mathcal{T}_{\ell_{1}}\cap\mathcal{T}_{\ell_{2}}|}{\min(|\mathcal{T}_{\ell_{1}}|,|\mathcal{T}_{\ell_{2}}|)}$
where $\mathcal{T}_{\ell}$ represents the set of items administered to model $\ell$.
C.5 Psychometric Interpretation Guidelines
Item Exposure Rate Benchmarks: Values below 20–30% are considered desirable in large-scale adaptive testing, as they indicate that no small subset of items dominates the evaluation, reducing security risks and promoting item bank longevity.
Test Overlap Rate Benchmarks: Values under 10–15% are typically considered good in adaptive testing practice, providing sufficient form diversity to resist contamination while maintaining measurement consistency.
Runtime Considerations: All measurements reflect computational efficiency in fully automated settings. Runtime includes item selection via Fisher information maximization but excludes model inference time, which varies significantly across architectures.
Appendix D Model Fit Evaluation across Datasets
Table 4 summarizes the average Root Mean Square Error of Approximation (RMSEA) from the MIRT M2 model fit across multiple benchmark datasets. Overall, all datasets exhibit low RMSEA values (ranging from 0.0436 to 0.0690), indicating a generally good fit between the latent multidimensional item response model and the observed response data. Among the evaluated datasets, GSM8K achieves the lowest average RMSEA (0.0436), suggesting that the underlying latent trait structure aligns particularly well with the response patterns, likely due to its consistent problem-solving format. In contrast, TruthfulQA shows a slightly higher RMSEA (0.0690), reflecting greater variability or multidimensionality in the response space.
The narrow gaps between the 5% and 95% RMSEA bounds across datasets indicate model stability and relatively low dispersion of fit indices among individual files. This consistency across diverse benchmarks such as Winogrande, HellaSwag, and ARC supports the reliability of the MIRT calibration process. Taken together, these results suggest that the applied MIRT M2 model provides a satisfactory and robust representation of item-level relationships across heterogeneous reasoning and language understanding tasks.
Table 4: Summary of MIRT M2 Average RMSEA across datasets
Dataset Avg. RMSEA Avg. 5% RMSEA Avg. 95% RMSEA Num. Files
Winogrande $0.0565$ $0.0561$ $0.0568$ $10$
TruthfulQA $0.0690$ $0.0686$ $0.0693$ 66
HellaSwag $0.0482$ $0.0478$ $0.0486$ $50$
GSM8K $0.0436$ $0.0433$ $0.0440$ $12$
ARC $0.0588$ $0.0585$ $0.0592$ 88
  • Note: ‘Avg. RMSEA‘ is the mean RMSEA across all files for each dataset. ‘Avg. 5% RMSEA‘ and ‘Avg. 95% RMSEA‘ are the averages of the file-level 5th and 95th percentile RMSEA bounds, respectively.
Appendix E Data Preprocessing Details
E.1 Point-Biserial Correlation Formula
The point-biserial correlation [23] for item $i$ is defined as:
$r_{pb}(i)\;=\;\frac{\bar{T}_{\,\ell\mid Y_{i\ell}=1}-\bar{T}_{\,\ell\mid Y_{i\ell}=0}}{s_{T}}\cdot\sqrt{p_{i}q_{i}},$
where $\bar{T}_{\,\ell\mid Y_{i\ell}=1}$ and $\bar{T}_{\,\ell\mid Y_{i\ell}=0}$ are the mean total scores of models that answered item $i$ correctly and incorrectly, respectively; $s_{T}$ is the standard deviation of total scores $\{T_{\ell}\}$; $p_{i}=\tfrac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}Y_{i\ell}$ is the proportion of models that answered item $i$ correctly; $q_{i}=1-p_{i}$; and $|\mathcal{L}|$ is the number of models.
E.2 Filtering Results
WinoGrande TruthfulQA HellaSwag GSM8K ARC
Num. Selected Models 4680 4635 3467 4195 4162
Num. Selected Item 1046 628 5608 1306 842
Avg. $M_{2}$ RMSEA 0.0565 0.0690 0.0482 0.0436 0.0588
Table 5: Summary of model and item filtering per benchmark. RMSEA values reflect average fit from $M_{2}$ statistics under the 3PL IRT model. Calibration and ability estimation were performed using the mirt package in R, with WLE as the estimator. Items were filtered based on variability, ceiling effects, and point-biserial discrimination.
Appendix F Evaluation Metric Definitions
This section provides the exact mathematical definitions of the evaluation metrics introduced in Section 4, along with brief interpretations.
Average Mean Absolute Error (MAE).
For each model $\ell$, let $\hat{\theta}_{\ell}$ denote the CAT-derived ability estimate and $\hat{\theta}_{\ell}^{\text{whole}}$ the full-bank reference estimate. The average MAE is:
$\text{MAE}=\frac{1}{|\mathcal{L}|}\sum_{\ell\in\mathcal{L}}\left|\hat{\theta}_{\ell}-\hat{\theta}_{\ell}^{\text{whole}}\right|,$ (4)
where $|\mathcal{L}|$ is the number of models evaluated. Interpretation: Lower MAE indicates higher fidelity, i.e., CAT ability estimates more closely match the whole-bank reference.
Average Item Exposure Rate.
Let $h_{i}$ denote the number of models administered item $i$, with $|\mathcal{I}|$ total items and $|\mathcal{L}|$ total models. The item exposure probability for item $i$ is
$P(A_{i})=\frac{h_{i}}{|\mathcal{L}|}.$ (5)
The average item exposure rate is then
$\bar{P}(A_{i})=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}P(A_{i}).$ (6)
Interpretation: Lower values indicate higher adaptivity and greater item diversity, while higher values suggest uniform or repetitive item usage across models.
Test Overlap Rate.
Following chen2005simplified [24], the expected proportion of common items between two randomly selected test forms is given by
$\bar{Q}=\frac{|\mathcal{L}|\sum_{i=1}^{|\mathcal{I}|}P(A_{i})^{2}}{\bar{L}(|\mathcal{L}|-1)}-\frac{1}{|\mathcal{L}|-1},$ (7)
where $\bar{L}$ is the average test length. Interpretation: Lower values of $\bar{Q}$ imply greater test form diversity, which reduces risks of collusion and item memorization.
Correlation Metrics
For completeness, we provide the definitions of the rank-based correlation coefficients used in Section 4.
Spearman correlation.
$\rho=1-\frac{6\sum_{i=1}^{n}d_{i}^{2}}{n(n^{2}-1)},$
where $d_{i}$ is the rank difference for observation $i$ across the two measures.
Kendall correlation.
$\tau=\frac{(\#\text{concordant pairs})-(\#\text{discordant pairs})}{\tfrac{1}{2}n(n-1)}.$
Appendix G Additional Experimental Results
While both TinyBenchmarks [8] and MetaBench [9] include MMLU [25] as part of their evaluation suites, they treat it as a single unified dataset by aggregating all 57 subject areas. In contrast, we perform evaluation on a per-subject basis. This design choice acknowledges the heterogeneous nature of MMLU, where each subject represents a distinct knowledge domain with varying linguistic characteristics, content distributions, and difficulty levels. Aggregating across all subjects can obscure these domain-specific patterns and limit interpretability in adaptive assessment.
Since each MMLU subject contains approximately 150 items, we randomly selected several representative subjects—Anatomy, Astronomy, and High-School Macroeconomics—to illustrate ATLAS’s behavior under data-sparse conditions. The corresponding results are reported in Table 6. Despite the small number of items per subject, ATLAS consistently demonstrates robust adaptive evaluation performance. As the selection threshold is relaxed ($SE\leq 0.1\rightarrow 0.3$), the mean absolute error (MAE) increases moderately (e.g., from 0.099 to 0.235 in Anatomy), while the number of evaluated items is substantially reduced (approximately 80%). This indicates that ATLAS effectively balances efficiency and accuracy, even in limited-data regimes.
Moreover, reductions in test overlap and exposure rates across the evaluated subjects suggest that the adaptive mechanism achieves broader item coverage and mitigates redundancy. Evaluation time also decreases proportionally with the number of items, confirming the computational efficiency of the adaptive process.
Table 6: Adaptive evaluation results on MMLU subsets (item range: 10–100). Lower is better $\downarrow$.
Benchmark Method MAE $\downarrow$ Avg. Item $\downarrow$ Test Overlap $\downarrow$ Exposure Rate $\downarrow$ Avg. Time (s) $\downarrow$
MMLU-Anatomy
(113 items)
ATLASSE≤0.1 0.099 100.00 0.94 0.88 4.93
ATLASSE≤0.2 0.149 53.29 0.54 0.47 2.70
ATLASSE≤0.3 0.235 20.49 0.32 0.18 0.98
MMLU-Astronomy
(136 items)
ATLASSE≤0.1 0.099 93.22 0.82 0.69 6.30
ATLASSE≤0.2 0.157 48.21 0.48 0.35 3.29
ATLASSE≤0.3 0.235 11.80 0.40 0.11 0.79
MMLU-High School
Macroeconomics (281 items)
ATLASSE≤0.1 0.089 87.94 0.48 0.31 10.49
ATLASSE≤0.2 0.197 27.64 0.22 0.13 3.46
ATLASSE≤0.3 0.240 18.26 0.20 0.09 2.34
Figure 4: Comparison of raw average scores and whole-bank ability estimates on TruthfulQA. (Left) While average scores compress performance at the extremes, whole-bank ability estimates reveal clearer separation among both low- and high-performing models, reflecting sensitivity to item difficulty and discrimination. (Right) Rank comparison shows strong consistency between the two measures (Spearman $\rho=0.97$, Kendall $\tau=0.87$), but ability-based ranking provides finer resolution, especially in distinguishing weaker and stronger models beyond what raw accuracy captures.
Figure 5: Comparison of raw average scores and whole-bank ability estimates on WinoGrande. (Left) Whole-bank estimates show a non-linear relationship with average score and reveal clearer separation on high-performing models, highlighting that ability captures relative item difficulty and provides finer differentiation beyond raw accuracy. (Right) Rank comparison indicates strong but imperfect alignment (Spearman $\rho=0.86$, Kendall $\tau=0.70$), with deviations from the diagonal reflecting cases where ability-based ranking distinguishes models more effectively than accuracy alone.
Figure 6: Comparison of raw average scores and whole-bank ability estimates on ARC. (Left) Whole-bank estimates exhibit a non-linear relationship with average scores, providing clearer separation on high-performing models by accounting for item difficulty and discrimination. (Right) Rank comparison shows strong but not perfect alignment between the two metrics (Spearman $\rho=0.91$, Kendall $\tau=0.77$), with deviations from the diagonal highlighting cases where ability-based ranking offers more informative distinctions than raw accuracy alone.
Figure 7: Two models with the similar average accuracy (0.713) and (0.714) on ARC nevertheless receive very different whole-bank ability estimates. Model A (mera-mix-4x7B) attains a whole-bank ability rank of 270 because its correct responses are concentrated on more difficult items. In contrast, Model B (LLaMAAntino-3-ANITA-8B-Inst-DPO-ITA) is assigned a much lower whole-bank ability rank of 2612, as its successes occur primarily on easier items. This divergence shows how IRT-based ability estimation can distinguish models that appear identical under raw accuracy by accounting for item difficulty.
Figure 8: Two models with the same average accuracy (0.853) on HellaSwag nevertheless receive very different whole-bank ability estimates. Model A (supermario_v1) attains a whole-bank ability rank of 347 because its correct responses are concentrated on more difficult items. In contrast, Model B (contaminated_proof_7b_v1.0_safetensor) is assigned a much lower whole-bank ability rank of 3074, as its successes occur primarily on easier items. This divergence shows how IRT-based ability estimation can distinguish models that appear identical under raw accuracy by accounting for item difficulty.